OpenStudy (anonymous):

Write an equivalent first-order differential equation and initial condition for y.

OpenStudy (anonymous):

\[y=-1\int\limits_{0}^{x}(2t-y(t))dt\]

OpenStudy (accessdenied):

Do you have any ideas so far?

OpenStudy (anonymous):

I'm as far as y'=

OpenStudy (anonymous):

I just need some direction on how to go about working this problem.

OpenStudy (anonymous):

What is first-order? I think I can do the integral.

OpenStudy (accessdenied):

First-order means the equation only contains the function and its first derivative.

OpenStudy (accessdenied):

So you are correct to take the derivative of both sides, the right-hand side should become a function of y and t and no extra higher order derivatives.

OpenStudy (anonymous):

derivative, or integral?

OpenStudy (accessdenied):

Here: \( \displaystyle y = - \int_{0}^{x} \left(2t - y(t)\right) \ dt \) If we integrated both sides, we would not have the first derivative anywhere. Instead, derivative of both sides: \( \displaystyle y' = \dfrac{d}{dx} \int_{0}^{x} \left(2t - y (t) \right) \ dt \) Note the derivative with respect to x because the right-hand integral is a function of x when it is evaluated.

OpenStudy (anonymous):

ahh, I see! Thanks!!

OpenStudy (anonymous):

derive both sides, and my function is on the right.

OpenStudy (accessdenied):

Yup. The right-hand side will end up leaving the function y(x). The d/dx of integral might be interesting to evaluate if you do not recall the first fundamental theorem of calculus, but if you treat integral of y(t) dt = Y(t) such that d/dt Y(t) = y(t), you can still evaluate by doing the integral first.

OpenStudy (anonymous):

y'(x) is f(x,y(x))

OpenStudy (anonymous):

so y'=2x-y

OpenStudy (anonymous):

and the initial condition is y(0)=-1

OpenStudy (anonymous):

Thanks for the help. I just needed some help jogging my memory.

OpenStudy (accessdenied):

that is about correct, I just saw I missed the negative sign in front when I wrote the derivative of both sides: \( \displaystyle y' = \color{red}{-} \frac{d}{dx} \int_{0}^{x} 2t - y(t) \ dt \) which was my bad. your right-hand side is just -1(2x - y) = y - 2x

OpenStudy (accessdenied):

although i don't see where you found y(0) = -1? although the -1 is multiplied on, we have an integral from 0 to 0 (covers no area).

OpenStudy (accessdenied):

just to emphasize the substitution: \( \displaystyle y \color{green}{(x)} = - \int_{0}^{\color{green}{x}} \left( 2t - y(t) \right) \ dt \) x=0: \( \displaystyle y(\color{red}{0}) = - \int_{0}^{\color{red}{0}} \left( 2t - y(t) \right) \ dt \)

OpenStudy (anonymous):

I'll have to get back to this later. I have a 3 year old climbing me! Thanks again.

OpenStudy (accessdenied):

glad to help! and good luck. :p in the end i do end up with y' = y - 2x and y(0) = 0. (I'll be getting off soon so I'll leave it here for any additional assistance).